OpenStudy (anonymous):

what is the radian measure of the smaller angle formed by the hands of a clock at 7 o'clock?

OpenStudy (anonymous):

150 degrees

OpenStudy (anonymous):

how'd u het that

OpenStudy (anonymous):

get

OpenStudy (anonymous):

@swagbear how did u gt that

OpenStudy (anonymous):

\[\frac{360^o}{12hours}=30^o/hr\] Relative to noon/midnight or when the larger "minute" hand is at that point \[7*30^o=210^o\ ] (or 360 minus this, depending on which way you want to look at it, which is 150 degrees)